- X sq + X + 1 = 0
What is X ?

Answer by sachi(548)

- X^2 + X + 1 = 0
or X^2 - X - 1 = 0
or 4X^2 -4 X -4 = 0 multiplying 4
or (2x)^2-2*2x*1+1-1-4=0
or (2x-1)^2=5
or 2x-1=+/-sqrt5
or x=(1+/-sqrt5)/2
ans